# CoAct-1: Computer-using Agents with Coding as Actions

### v1

@@ 1 Introduction
Our experimental analysis provides strong evidence for the advantages of this hybrid design. On the comprehensive OSWorld benchmark, CoAct-1 establishes a new state-of-the-art, achieving an overall success rate of 56.17%. This marks a significant improvement over leading baselines like GTA-1 (45.20%). The performance gains are most pronounced in categories where programmatic control is highly advantageous. For instance, in OS-level (79.16%), multi-application (43.73%), and Thunderbird email (80.00%) tasks, the Programmer agent’s ability to execute precise scripts leads to substantial gains over the strongest GUI-only methods. Beyond improving success rates, our dual-modality approach dramatically enhances operational efficiency. By replacing long, error-prone click sequences with concise code, CoAct-1 solves tasks in an average of just 10.54 steps, a stark contrast to the 15 steps required by agents like GTA-1. This efficiency underscores the potential of our approach to pave a more robust and scalable path toward generalized computer automation.
@@ 4.3 Implementation Details
CoAct-1 settings We implement CoAct-1 using AG2 (Wu et al., 2023). We tried different backbones for the Programmer and Orchestrator, including OpenAI o3 (OpenAI, 2025b) and o4-mini (OpenAI, 2025b). For the GUI Operator, we use computer-use-preview, a vision-language action model finetuned by OpenAI for computer use, as the backbone model. We use the o4-mini as the summarizer for summarizing the conversation history between the Programmer and the Orchestrator. We set the max round $I$ for Programmer as 20, max step $K$ for GUI Operator as 25, and max round $J$ for Orchestrator as 15. Therefore, the number of system interactions, i.e, the number of steps, for CoAct-1 is upper bounded by 375. For Table 2
@@ 4.5 Discussion
The analysis of CoAct-1’s operational efficiency, illustrated in Figure 3, reveals that our hybrid approach is substantially more efficient than leading GUI-only agents. This efficiency is a key factor in its improved success rate. As shown in 3(a), CoAct-1 solves tasks with an average of 10.15 steps. This represents a significant improvement over other high-performing agents like GTA-1, which requires 15.22 steps, and UI-TARS, which needs 14.90 steps on average. While OpenAI CUA 4o averages fewer steps (6.14) , its overall success rate is much lower compared to CoAct-1’s (31.40% v.s. 59.93%). This indicates that CoAct-1’s efficiency is coupled with greater effectiveness. The source of this efficiency lies in the strategic use of coding actions. 3(b) supports this by showing that coding actions help keep the total steps per task relatively low. This efficiency is crucial for robust performance. 3(c) shows that coding is particularly beneficial in complex domains like "LibreOffice Calc", "Multi-apps" and direct OS interactions, where a large proportion of tasks are solved with code. A single script can replace a long and error-prone sequence of GUI clicks, streamlining the workflow. 3(d) illustrates a clear trend: tasks that require more actions are more likely to fail. By reducing the total number of steps, the hybrid approach not only accelerates task completion but also minimizes the opportunities for error. The ability to dynamically select the most appropriate action—either a direct coding command or a GUI interaction—is fundamental to the enhanced efficiency and reliability of CoAct-1.

### v2

@@ 1 Introduction
Our experimental analysis provides strong evidence for the advantages of this hybrid design. On the comprehensive OSWorld benchmark, CoAct-1 establishes a new state-of-the-art, achieving an overall success rate of 60.76%. This marks a significant improvement over leading baselines like GTA-1 (53.10%). The performance gains are most pronounced in categories where programmatic control is highly advantageous. For instance, in Calc (70.21%), multi-application (47.88%), and VS Code (78.26%) tasks, the Programmer agent’s ability to execute precise scripts leads to substantial gains over the strongest GUI-only methods. Beyond improving success rates, our dual-modality approach dramatically enhances operational efficiency. By replacing long, error-prone click sequences with concise code, CoAct-1 solves tasks in an average of just 10.15 steps, a stark contrast to the 15 steps required by agents like GTA-1. This efficiency underscores the potential of our approach to pave a more robust and scalable path toward generalized computer automation.
@@ 4.3 Implementation Details
CoAct-1 settings We implement CoAct-1 using AG2 (Wu et al., 2023). We tried different backbones for the Programmer and Orchestrator, including OpenAI o3 (OpenAI, 2025b) and o4-mini (OpenAI, 2025b). For the GUI Operator, we use computer-use-preview, a vision-language action model finetuned by OpenAI for computer use, as the backbone model. We use the o4-mini as the summarizer for summarizing the conversation history between the Programmer and the Orchestrator. We set the max round $I$ for Programmer as 20, max step $K$ for GUI Operator as 25, and max round $J$ for Orchestrator as 15. Therefore, the number of system interactions, i.e, the number of steps, for CoAct-1 is upper bounded by 375.
@@ 4.5 Discussion
The analysis of CoAct-1’s operational efficiency, illustrated in Figure 3, reveals that our hybrid approach is substantially more efficient than leading GUI-only agents. This efficiency is a key factor in its improved success rate. As shown in 3(a), CoAct-1 solves tasks with an average of 10.15 steps. This represents a significant improvement over other high-performing agents like GTA-1, which requires 15.22 steps, and UI-TARS, which needs 14.90 steps on average. While OpenAI CUA 4o averages fewer steps (6.14) , its overall success rate is much lower compared to CoAct-1’s (31.40% v.s. 59.93% on 100 steps). This indicates that CoAct-1’s efficiency is coupled with greater effectiveness. The source of this efficiency lies in the strategic use of coding actions. 3(b) supports this by showing that coding actions help keep the total steps per task relatively low. This efficiency is crucial for robust performance. 3(c) shows that coding is particularly beneficial in complex domains like "LibreOffice Calc", "Multi-apps" and direct OS interactions, where a large proportion of tasks are solved with code. A single script can replace a long and error-prone sequence of GUI clicks, streamlining the workflow. 3(d) illustrates a clear trend: tasks that require more actions are more likely to fail. By reducing the total number of steps, the hybrid approach not only accelerates task completion but also minimizes the opportunities for error. The ability to dynamically select the most appropriate action—either a direct coding command or a GUI interaction—is fundamental to the enhanced efficiency and reliability of CoAct-1.